# Think Twice Before You Judge: Mixture of Dual Reasoning Experts for Multimodal Sarcasm Detection

### v1

@@ 7.1. Need of External Rationale
External reasoning is crucial in sarcasm detection, particularly when textual and visual cues alone fail to establish contradiction. For instance, in Figure 5(a), the image shows Easter-themed products on a shelf, while the text questions when Halloween products will arrive, implying that Easter items are being sold too early. A model without external knowledge may struggle to detect sarcasm since the image-text itself does not explicitly indicate any contradiction. However, external reasoning clarifies that Easter eggs appearing in January is unusual, reinforcing the sarcasm. MiDRE successfully classifies this instance as sarcastic, with a high gate value for External Rationale (ER) module (0.876), while the baseline and MiDRE without ER fail. In Figure 5(b), the text references a famous sports trade (Mike Ditka’s excessive trade for Ricky Williams) to suggest that the Broncos should do something similarly extreme. The accompanying image with confused expressions enhances the sarcasm. Without knowledge of the Ditka trade, the sarcasm might be missed, as the text could be interpreted as a genuine suggestion. External reasoning bridges this gap by explaining the historical context, leading MiDRE to detect sarcasm correctly (ER gate value: 0.921). In the case of Figure 5(c), both the image and text lack strong sarcastic elements. The image depicts a casual mirror selfie, and the text states that the person probably takes selfies with filters. MiDRE assigns a high ER gate value (0.781), suggesting that the Internal Reasoning (IR) module might be uncertain due to the overtly positive tone of the text. However, ER helps by recognizing that there is no cultural or contextual contradiction that would indicate sarcasm.
@@ 7.2. Impact of Noisy Rationale
External rationales, even when noisy or indirect, enhance the internal reasoning (IR) module by providing valuable contextual cues for sarcasm detection. For example, in Figure 7(a), MiDRE without external reasoning (ER) fails to detect sarcasm, while MiDRE succeeds by leveraging the image rationale, which highlights the neutral design of the seat. This neutrality amplifies the exaggerated praise in the text, making the sarcasm more apparent. Similarly, in 7(b), although the image rationale doesn’t directly contradict the caption, it describes the dog’s playful behavior, subtly clashing with the claim of it being ”well-behaved” at the vet. In both cases, ER offers factual descriptions that help IR identify subtle incongruities, demonstrating the complementary role of external rationales in improving sarcastic intent recognition.
@@ 8. Error Analysis
Our model, MiDRE, fails in cases where both internal and external reasoning fall short. In Figure 8(a), the image explicitly mentions ”sarcasm,” leading to model bias—falsely associating the term with sarcasm without deeper analysis. In Figure 8(b), sarcasm emerges from the contradiction of the makeshift arrangement of sticking a calendar to the wall using a tape with the caption, “I’m really classy.” Here, internal reasoning struggles to detect the subtle contradiction, while external reasoning lacks the context to flag it, causing the model to miss the sarcasm.
@@ 9. Conclusion
10. Appendix
10.1. Detailed Prompts
This section outlines the prompts in full details that we use for the rationale generation module.
10.1.1. Image Prompt
You are given an image. Follow the steps below:### Task 1: Image Analysis- Carefully analyze the image, considering any visible elements, context, and any potential irony or contrast in the visual content.- Based on your analysis, determine if the image contains sarcasm.  - Print **1** if the image is sarcastic.  - Print **0** if the image is not sarcastic.### Task 2: Reasoning- Provide a **concise** reasoning for why you chose the label in Task 1.- Ensure that your reasoning explains any ironic or sarcastic elements or the lack thereof in the image.### Output Format:- **Label**: [1/0]- **Reason**: [<reason>]### Notes:- The **Label** is determined by whether you believe the image is sarcastic or not.- The **Reason** should be a **concise** explanation directly related to the content of the image.- The tasks should be handled **independently**, but logically connected (task 2 explains task 1).- Your output should **strictly follow** this format without any additional commentary or explanation:# Adhere strictly to the order and independence of tasks as described. Provide outputs exactly in the specified format.
10.1.2. Text Prompt
You are given a text. Follow the steps below:### Task 1: Image Analysis- Carefully analyze the text, considering any exaggeration, incongruity or ironic elements in the textual content.- Based on your analysis, determine if the text contains sarcasm.  - Print **1** if the text is sarcastic.  - Print **0** if the text is not sarcastic.### Task 2: Reasoning- Provide a **concise** reasoning for why you chose the label in Task 1.- Ensure that your reasoning explains any ironic or sarcastic elements or the lack thereof in the text.### Output Format:- **Label**: [1/0]- **Reason**: [<reason>]### Notes:- The **Label** is determined by whether you believe the text is sarcastic or not.- The **Reason** should be a **concise** explanation directly related to the content of the text.- The tasks should be handled **independently**, but logically connected (task 2 explains task 1).- Your output should **strictly follow** this format without any additional commentary or explanation:# Adhere strictly to the order and independence of tasks as described. Provide outputs exactly in the specified format.
10.1.3. Multimodal Prompt
You are provided with an **Image** and a **Supporting Text**. Perform the following tasks independently:# **Definitions**:# - **Supporting Text**: The text provided separately, not embedded in the image.# - **Caption**: The text embedded within the image.1. **Extract Caption**:   - Extract the caption (if any) from the image. This step should not involve analyzing sarcasm.2. **Analyze Supporting Text**:   - Consider only the provided supporting text and ignore the caption or any information from the image.   - Is the supporting text sarcastic? Answer `1` for sarcastic and `0` for non-sarcastic.   - Provide reasoning based only on the supporting text.3. **Analyze Image and Caption**:   - Analyze the image and its caption together, considering any contrasts, contradictions, or ironic elements.   - Is the image-caption pair sarcastic? Answer `1` for sarcastic and `0` for non-sarcastic.   - Provide reasoning based on the image-caption pair.**Output Format**:Caption in Image: [Caption]Supporting Text Analysis:  - Label: [1/0]  - Reason: [<reason>]Image-Caption Analysis:  - Label: [1/0]  - Reason: [<reason>]**Guidelines**:- Keep the analysis of the supporting text completely independent of the image and caption.- Provide outputs in the specified format without additional commentary.- Clearly separate reasoning for the supporting text and the image-caption pair.
10.2. Performance of Reverse Configuration
Table 5. Performance on the MMSD2.0 dataset.
Model Acc. P R F1 MiDRE (w reverse) 82.13 81.62 80.20 80.90

### v2

@@ 7.3. Need of External Rationale
External reasoning is crucial in sarcasm detection, particularly when textual and visual cues alone fail to establish contradiction. For instance, in Figure 6(a), the image shows Easter-themed products on a shelf, while the text questions when Halloween products will arrive, implying that Easter items are being sold too early. A model without external knowledge may struggle to detect sarcasm since the image-text itself does not explicitly indicate any contradiction. However, external reasoning clarifies that Easter eggs appearing in January is unusual, reinforcing the sarcasm. MiDRE successfully classifies this instance as sarcastic, with a high gate value for ERE module (0.876), while the baseline and MiDRE without ERE fail. In Figure 6(b), the text references a famous sports trade (Mike Ditka’s excessive trade for Ricky Williams) to suggest that the Broncos should do something similarly extreme. The accompanying image with confused expressions enhances the sarcasm. Without knowledge of the Ditka trade, the sarcasm might be missed, as the text could be interpreted as a genuine suggestion. External reasoning bridges this gap by explaining the historical context, leading MiDRE to detect sarcasm correctly (ERE gate value: 0.921). In the case of Figure 6(c), both the image and text lack strong sarcastic elements. The image depicts a casual mirror selfie, and the text states that the person probably takes selfies with filters. MiDRE assigns a high ERE gate value (0.781), suggesting that the IRE module might be uncertain due to the overtly positive tone of the text. However, ERE helps by recognizing that there is no cultural or contextual contradiction that would indicate sarcasm.
@@ 7.4. Impact of Noisy Rationale
External rationales, even when noisy or indirect, enhance the IRE module by providing valuable contextual cues for sarcasm detection. For example, in Figure 9(a), MiDRE without ERE fails to detect sarcasm, while MiDRE succeeds by leveraging the image rationale, which highlights the neutral design of the seat. This neutrality amplifies the exaggerated praise in the text, making the sarcasm more apparent. Similarly, in 9(b), although the image rationale doesn’t directly contradict the caption, it describes the dog’s playful behavior, subtly clashing with the claim of it being ”well-behaved” at the vet. In both cases, rationales offers factual descriptions that help IRE identify subtle incongruities, demonstrating the complementary role of external rationales in improving sarcastic intent recognition.
@@ 8. Error Analysis
Our model, MiDRE, fails in scenarios where both the ERE and IRE produce weak or misleading signals. In Figure 10(a), the ERE gate value (0.413) is lower than the IRE gate value (0.587), indicating heavier reliance on internal reasoning. However, the internal reasoning is biased by the explicit presence of the word “sarcasm” in the image text, leading to an incorrect prediction. In Figure 10(b), the ERE gate value (0.629) outweighs the IRE gate value (0.371), suggesting the model leaned more on external reasoning. Here, the humor comes from the visual-textual contradiction between the caption “I’m really classy” and the taped wall calendar. The internal reasoning misses this subtle contradiction, while external reasoning fails to connect the visual setup with its implied sarcasm. Consequently, the model predicts this post as not sarcasm.
@@ 9. Conclusion
Ethical Considerations
This work on multimodal sarcasm detection is conducted entirely on publicly available datasets, with all private, personally identifiable content removed during preprocessing to ensure compliance with data protection and ethical research guidelines. The developed model is intended solely for academic and research purposes, and is not optimized, validated, or certified for use in production systems. While the model can provide insights into sarcasm patterns in multimodal data, its outputs should be interpreted with caution, as sarcasm detection is inherently subjective and may be influenced by cultural, contextual, and individual differences. We strongly discourage the use of this model in high-stakes or sensitive decision-making contexts, including but not limited to legal, political, or law enforcement scenarios.